OpenStudy (anonymous):

simple trig identity! prove sinx/1+cosx= 1/sinx - 1/tan x

OpenStudy (anonymous):

so i think you write not right equal

OpenStudy (anonymous):

you should do this on your own. it will help you learn about manipulation of trigonometric identities.

OpenStudy (anonymous):

honestly, i would if i knew how. i have 2 assignments of an online functions course left to do and once i'm done, i never have to do math again, so all i really need to do is finish them. i've done what i can on my own and now i'm asking for help.

OpenStudy (anonymous):

Note \[ 1-\cos^2x = \sin^2x \] And think about what you can multiply the top and bottom of the LHS (or RHS) to use this fact.

OpenStudy (anonymous):

tan(x)= sin(x)/cos(x)

OpenStudy (anonymous):

My method is more interesting ¬_¬

OpenStudy (anonymous):

\[a^2 - b^2 = (a+b)(a-b) \] May also be useful.

OpenStudy (anonymous):

I like Newton's method

OpenStudy (anonymous):

this is all just confusing me more. it's so much easier to understand when i see a solution.

OpenStudy (anonymous):

Sorry, I can't do your problem for you. Ask one of the other people.

OpenStudy (anonymous):

so far i took the right side and got 1/sinx-cosx/sinx which = 1-cosx/sinx but i dont know if that is correct, or what to do next

OpenStudy (anonymous):

That is correct, and useful. How can you get from that to the LHS?

OpenStudy (anonymous):

that's what i don't know.

OpenStudy (anonymous):

i some how need to flip it and chance the - to a +

OpenStudy (anonymous):

Not technically true. Because on the LHS, the sin(X) is on the top, but on the RHS, the sin(x) is on the bottom. See my above advice.

OpenStudy (anonymous):

but there are no ^2 in this proof

OpenStudy (anonymous):

OK. So if you take your LHS and RHS in the new form you have. \[\frac{\sin x}{1+\cos x} = \frac{1-\cos x}{\sin x} \] Multiplying both sides by (1+cosx)(sinx) gives a result that is clearly true. (s^2 - 1-c^2) HOWEVER it is best to go directly from LHS -> RHS (or vice versa) and not assume the result and work from that. So can you see how something similar could get from one side to the otehr directly?